OpenStudy (anonymous):

how am i suppose to determine the domain and range for the function: y= the square root of (x-4)? i dont want the answer i just want to understand the problem

OpenStudy (amistre64):

domain is all usable values for "x" the range is all the results for "y" that you get by using the elements of the domain

OpenStudy (amistre64):

ask yourself, does the square root function have any restrictions? what makes it go bad?

OpenStudy (anonymous):

wait so the domain is all real numbers? why?

OpenStudy (amistre64):

if the range and domain of the inverse is easier to find; that would be another method

OpenStudy (amistre64):

does the square root function have an restrictions? for example sqrt(-4) = ?? sqrt(0) = ?? sqrt(9) = ?? can we solve these?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

we cant solve any of them?

OpenStudy (anonymous):

well isnt the sqr of -4 an imaginary number?

OpenStudy (anonymous):

or something?

OpenStudy (amistre64):

yes it is, if we are allowed to step out of the reals

Parth (parthkohli):

The question is not very well worded...:| It's supposed to say that we have a "Real function".

OpenStudy (amistre64):

if your allowed to use the set of Complex numbers, then the domain is all complex values

OpenStudy (anonymous):

@ParthKohli yeah thats right

OpenStudy (amistre64):

or rather, if y is allowed to be an imaginary number; then the domain can include the set of reals

OpenStudy (amistre64):

\[\sqrt{u(x)}~:~u(x)=x-4\] \[u(x)\ge0\] \[x-4\ge0\] will define our domain if we assume sqrt to be a real valued function

OpenStudy (anonymous):

wait im confused

OpenStudy (amistre64):

youll have to be more specific about your confusion

OpenStudy (anonymous):

i want to know something concrete about how im suppose to solve problems like this kind of like a rule or certain steps because i dont understand the concept

OpenStudy (amistre64):

domain is the set of elements that are allowable: lets say that you have 20 dollars to spend in the store; your domain is then $20 the range of options for you to purchase is then within the values of free to $20 less taxes

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

the steps for the mathical stuff is: determine usable values for the given function determine the values of "x" that give us those values if the domain has a least value, plug that in to the least value of the range if the domain has a highest value, plug that in to the highest value of the range

OpenStudy (amistre64):

so, what values, if any, do we have to omit from the square root function for it to be give us real values?

OpenStudy (anonymous):

isnt there alot of options to plug in?

OpenStudy (amistre64):

i spose we could name them one by one, or it might be easier to group them into a specific set that defines what they all have in common

OpenStudy (anonymous):

oh so like all real numbers? does that go for the range too?

OpenStudy (amistre64):

all real numbers is not correct

OpenStudy (amistre64):

sqrt(-4) does not produce a real number does it?

OpenStudy (anonymous):

ok will it be all real numbers greater than -4?

OpenStudy (amistre64):

we might be thinking of 2 different things here, im not focused on the (x-4) part just yet im just trying to define the values that either make sqrt a real function, or not

OpenStudy (amistre64):

sqrt(-3) is not a real value; sqrt(-2) is not a real value; sqrt of any negative number does not produce a real value

OpenStudy (amistre64):

negative number are less than 0 so, when is x-4 < 0 ? or to say the same thing differently when is x-4 >= 0 ?

OpenStudy (amistre64):

teh second setup will be more useful to us to define the domain with since it produces all the values of x that we can actually use

OpenStudy (anonymous):

do you mean greater than or equal to zero for the second setup ? so i guess 4 is the lowest number that can be plug into the second setup for it to equal 0

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so all numbers greater than or equal to four?

OpenStudy (amistre64):

that is our domain, yes

OpenStudy (amistre64):

the range is found by using the domain at x=4; y=0 at x=5; y=1 at x=8, y=2 at x=13; y=3 without getting into calculus, its hard to determine if there is a maximum value for y; but i already know that \(\sqrt{\infty-4}=\infty\)

OpenStudy (amistre64):

\[y=\sqrt{x-4}\] \[y'=\frac{1}{2\sqrt{x-4}}\] since y' can never be zero in this case, there is no maximum, and the function is always increaseing

OpenStudy (amistre64):

gotta get to class, good luck ;)

OpenStudy (anonymous):

ok thanks for all your time and help